OpenStudy (anonymous):

Is this equation linear or nonlinear? 5xy + 4 = 9

OpenStudy (anonymous):

The expression is, 5xy+4=9 =>5xy=9-4=5 =>xy=5/5=1 =>y=1/x which is a equation of rectangular hyperbola and thus nonlinear equation.